OpenStudy (anonymous):

Divide and Simplify: y^2-81over 64y+576 divide y-9 over 72 really need help with this one please show steps.

OpenStudy (anonymous):

\[\frac{y^2-81}{64y+576}\]\[=\frac{(y+9)(y-9)}{64(y+9)}\]\[=\frac{y-9}{64}\]I don't understand the last part of your question above.

OpenStudy (anonymous):

that is how it is y^2-81 over 64y+576 that is one section then divide y-9 over72

OpenStudy (anonymous):

so divide my answer by \[\frac{y-9}{72}\]???

OpenStudy (anonymous):

no its all one equation y^2 over 64y+576/y-9 over 72 does that help any

OpenStudy (anonymous):

No, that doesn't help. Can you use parentheses to clarify?

OpenStudy (anonymous):

or the equation editor?

OpenStudy (anonymous):

(y^2-81 over 64y+576)/(y-9 over 72)

OpenStudy (anonymous):

its in fraction form

OpenStudy (anonymous):

when I say over

OpenStudy (anonymous):

so yes, divide my original answer by \[\frac{y-9}{72}\]which is eqiivalent to multiplying by the reciprocal:\[\frac{y-9}{64}\times\frac{72}{y-9}\]the y-9's cancel, leaving you with 72/64, or 9/8.

OpenStudy (anonymous):

Thank you